OpenStudy (anonymous):

A random sample of 130 students is chosen from a population of 4,500 students. If the mean IQ in the sample is 120 with a standard deviation of 5, what is the 99% confidence interval for the students' mean IQ score? 118.87−121.13 107.12−132.88 125−135 115−125

OpenStudy (anonymous):

dam you smart

OpenStudy (anonymous):

now you have reached the outer limits of my abilities good luck with random math class

OpenStudy (anonymous):

xD its a mix of different things. this im not so sure.

OpenStudy (anonymous):

this i have no clues

OpenStudy (aum):

Look up the z-scores table for 99% confidence interval. It is 2.576. The 99% confidence interval = Mean - 2.576 * SD to Mean + 2.576 * SD where SD is standard deviation. The 99% confidence interval = 120 - 2.576 * 5 to 120 + 2.576 * 5 = 107.12 to 132.88